A students is going to graduate from a business department of a university by the end of the
semester. After being interviewed at two companies he likes, he assesses that his probability of
getting an offer from company A is 0.6, and his probability of getting an offer from company B is
0.45. If he believes that the probability that he will get offers from both companies is 0.35, what
is the probability that he will get at least one offer from these two companies?
1
Expert's answer
2020-06-10T19:11:15-0400

Let A be the event that the student will get an offer from company A

and B be the event that the student will get an offer from company B

Then A "\cap" B is the event that he will get offers from both companies


We are given,


P(A) = 0.6, P(B) = 0.45 and P(A "\cap" B) = 0.35


"\therefore" The probability that he will get at least one offer from these two companies is


P(A "\cup" B) = P(A) + P(B) - P(A "\cap" B) = 0.6 + 0.45 - 0.35 = 0.7


Answer: The probability that the student will get at least one offer from these two companies is 0.7.